OpenStudy (av0cad0):

2/3 (3x + 9) = -2 (2x + 6) I just need a walkthrough (my teacher isn't that great with this review)

OpenStudy (av0cad0):

@zepdrix @mathmate

OpenStudy (mathmate):

Is this \(\frac{2}{3}(3x+9)=-2(2x+6)\) ? What would be your first step?

OpenStudy (av0cad0):

Yes, it is.

OpenStudy (mathmate):

So would you please show me what your first step would be to solve this equation?

OpenStudy (av0cad0):

Thats the thing: I don't know. I've been taking notes in class that are all algebra and how to do it, but there is litterally 5 questions that look like that on this homework sheet.

OpenStudy (mathmate):

The problem is not with your teacher. It's with you, unfortunately. Before you go to the step of solving equations, you should have learned and mastered the distributive properties of algebraic equations, namely how to "remove parentheses" from expressions. If you had learned and understood the previous steps ("removing parentheses"), you will have no problem proceeding to solve equations. For example, to solve 4(x+2)=2(x-3) the first step (and reflex) is to use distributive properties to "remove the parentheses", 4x+8=2x-6 Then isolate the unknown (x) on the left side, and known (numbers) on the right, 4x-2x=-6-8 Proceed to simplify 2x=-14 divide by the coefficient of the unknown (2) to get the final answer: x=-14/2=-7 Follow the same steps and see if you can work out your example. I did not use your question in order not to deprive you of practice, which is what you need. Post what you have, and if in doubt, post what you've got.

jimthompson5910 (jim_thompson5910):

With me, fractions always complicate things. When we have them in equations, we can multiply both sides by the LCD to clear out the fractions. In this case, there's only one fraction 2/3. The denominator here is 3. So multiply both sides by 3 to clear out that fraction \[\Large \frac{2}{3}(3x+9) = -2(2x+6)\] \[\Large {\color{red}{3*}}\frac{2}{3}(3x+9) = {\color{red}{3*}}(-2)(2x+6)\] \[\Large 3x+9 = -6(2x+6)\] Hopefully you're able to see how 3 times 2/3 equals 2. The '3's cancel out here.

OpenStudy (av0cad0):

@mathlete If you're going to insult me, i'm not taking any of your "help". There is nothing "wrong" with me, you don't even know who I am. If anything i'm just another poster to you. I take 3/4 Pre-AP classes, and if I say "my teacher isn't helping", then my teacher isn't helping. Don't post in any of my future questions, please. I by no means need or want help from someone who is rude.

OpenStudy (av0cad0):

Also before I forgot, Jim that is all I needed. Thank you.